in a double displacement creation 15.84g of ammonium phosphate completely reacts with excess lead (iv) nitrate, determine the mass of the precipitate formed when these 2 solutions are mixed


1
Expert's answer
2020-10-23T09:15:26-0400

4(NH4)3PO4(aq) + 3Pb(NO3)4(aq) → Pb3(PO4)4(s) + 12NH4NO3(aq)

M((NH4)3PO4) = 149.08 g/mol

n = m/M

n((NH4)3PO4) "= \frac{15.84}{149.08} = 0.106 \;mol"

n(Pb3(PO4)4) = 1/4n((NH4)3PO4) "= 0.25 \times 0.106 = 0.0265 \;mol"

"m = n \times M"

M(Pb3(PO4)4) = 1001.48 g/mol

m(Pb3(PO4)4) "= 0.0265 \times 1001.48 = 26.54 \;g"

Answer: 26.54 g